OpenStudy (anonymous):

please help me 3u-4v=9 5u-8v=4

OpenStudy (amistre64):

\[{{3\ 9}\choose{5\ 4}}^{(45)}_{12}=-45+12=-33\] \[{{3\ -4}\choose{5\ -8}}^{(-20)}_{-18}=20-18=2\] v = -33/2 if i did that right

OpenStudy (anonymous):

what is this method?

OpenStudy (anonymous):

its systems of equations

OpenStudy (amistre64):

something with determiates and Dv/D i think

OpenStudy (anonymous):

algebra2

OpenStudy (anonymous):

the answeres are (14,33/4) i just dont know how to get that

OpenStudy (amistre64):

33/4 ?

OpenStudy (anonymous):

yeah its a fraction

OpenStudy (mathmate):

The second determinant is -4

OpenStudy (amistre64):

8*3 = 24 .. not 18 lol

OpenStudy (anonymous):

is this some of your variation of cramer's rule ?

OpenStudy (amistre64):

its something like that; but I spent the winter break reading a linear algebra book to prep me for class

OpenStudy (anonymous):

yes they are sets of equations but is that suppose to be the answere?

OpenStudy (amistre64):

\[{{-4\ 9}\choose{-8\ 4}}^{(-72)}_{-16}=-16+72=56\] and for this one we just negate the determinate for the denominator 56/4 = 14

OpenStudy (anonymous):

lol, spare the kid :P

OpenStudy (amistre64):

;) -2(3u-4v=9) 5u-8v=4 -6u+8v=-18 5u-8v= 4 ------------ -u = -14 u = 14

OpenStudy (anonymous):

btw is this approach optimal for more equations? I mean this seems a bit convoluted for \( 2\times 2 \).

OpenStudy (amistre64):

linear algebra is never optimal ...

OpenStudy (anonymous):

hm I guess that's why we have numerical analysis ... ;)

OpenStudy (amistre64):

\[{{3\ -4}\choose{5\ -8}}{u\choose v}={9\choose 4}\] \[{u\choose v}={{3\ -4}\choose{5\ -8}}^{-1}{9\choose 4}\]

OpenStudy (anonymous):

the one were you put the winky face was only the first 1 right ???

OpenStudy (amistre64):

\[{{3\ -4}\choose{5\ -8}}^{-1}=\frac{1}{det}{{-8\ 4}\choose{-5\ 3}}\] \[{{3\ -4}\choose{5\ -8}}^{-1}=\frac{1}{-4}{{-8\ 4}\choose{-5\ 3}}\] \[{{3\ -4}\choose{5\ -8}}^{-1}={{2\ -1}\choose{5/4\ -3/4}}\] such fun

OpenStudy (anonymous):

When you got u= 14 you can put 42-4x=9 and solve for x.

OpenStudy (anonymous):

should get 33/4

OpenStudy (anonymous):

thanks bro ! really helped

OpenStudy (anonymous):

i tried solving that last part but i got 42/13 ???

OpenStudy (amistre64):

\[{u\choose v}={{2\ -1}\choose{5/4\ -3/4}}{9\choose 4}\] \[{u\choose v}=9{2\choose5/4}^{18}_{45/4} +4{-1\choose -3/4}^{-4}_{-12/4}={{14}\choose{33/4}}\]

OpenStudy (amistre64):

.... i needed the practice :)

OpenStudy (anonymous):